# MAAdvisor: Zero-Shot Index Advisor using Multi-Agent LLMs

### v1

@@ 6.3. Multi-Agent Pipeline
Once the global and local agents are developed, they can be integrated into a hierarchical multi-agent pipeline, enabling zero-shot scalable index recommendation. As shown in Algorithm 1, the multi-agent pipeline receives the storage budget and the column candidates as inputs. To balance efficiency with effectiveness, we introduce a hyperparameter , which defines the maximum number of planning sub-steps. At the beginning of each sub-step, the planning agent determines the next action. If additional indexes can be selected, the selection agent chooses a single-column index from the effectively represented column candidates. When combining existing indexes into a composite index could further improve workload performance, the combination agent refines the current recommendation by composing these indexes. Since the selected indexes may contain duplicates, regressions, or low-impact choices, the revision agent ensures effectiveness by removing such cases. After each action, the reflection agent reviews prior decisions and guides subsequent planning. Through the planning agent’s reasoning, if the index configuration is deemed optimal, iterations terminate early, and the final selection is output as the index recommendation.
7. Experiments
7.1. Experimental Setup
7.1.1. Dataset.
We select four different datasets from both hypothetical and real-world settings for our evaluations.
TPC-H: The TPC-H dataset is a widely used benchmarking dataset for measuring the performance of database management systems. We generate the data with a scale of 10 (around 10 GB). The TPC-H benchmark has in total of 22 different query templates. In order to simplify the analysis and make a fair comparison of the methods, we follow previous methods (Siddiqui et al., 2022; Schlosser et al., 2019) to remove the query templates 2, 17, 20 that contribute significantly to the total execution time of the workload or have little potential improvement from indexes. The test workload for TPC-H has in total of 19 queries randomly generated from the 19 query templates left.
TPC-DS: The TPC-DS dataset is another widely used benchmark with a more complicated database schema and query templates. Data with a scale of 10 (around 10 GB) is generated. The TPC-DS benchmark has 99 different query templates, and for the same reason in the TPC-H benchmark, we remove the query templates 4, 6, 9, 10, 11, 14, 23, 24, 32, 35, 36, 39, 41, 70, 86, and 95. The test workload for TPC-DS consists of 83 queries generated from the 83 query templates left.
DSB: The DSB benchmark (Ding et al., 2021) is derived from the TPC-DS benchmark, while enhancing the TPC-DS with complex data distribution and challenging yet semantically meaningful query templates. DSB has in total of 15 templates for single block queries, 22 templates for multi-block queries, and another 15 templates for SPJ queries, especially evaluating techniques with limited capabilities. In our experiments, we only utilized the general-purpose 37 query templates in the first two groups and constructed a workload of 37 queries randomly generated from the templates.
IMDB (JOB workload): The IMDB dataset (Maas et al., 2011) is a collection of information about movies, TV shows, and actors. Join order benchmark (JOB (Leis et al., 2015)) is an OLAP benchmark querying the IMDB dataset and designed to assess the capabilities of a database management system (DBMS) in optimizing and executing complex real-world join queries. The JOB workload has 113 real-world queries.
7.1.2. Experimental Environments.
We choose the popular DBMS PostgreSQL 14.2 to evaluate all the queries (Pos, [n.d.]). We follow the official documentaries and websites (TPC, [n.d.]; Leis et al., 2015; dsb, [n.d.]) to create PostgreSQL databases for all the above datasets. All the experiments are conducted on an isolated server with two Intel(R) Xeon(R) Platinum 8469C CPUs featuring 46 cores and 96 threads, along with 460GB of RAM.
7.1.3. Baseline Methods.
Heuristic-based baselines: Heuristic-based index advisors can be broadly categorized into two types by their optimization approaches. Those following an ADD strategy incrementally build a set of indexes by selecting candidates one by one. In our experiments, we include Anytime (Chaudhuri and Narasayya, 2020), DB2Advis (Skelley, 2000), Extend (Schlosser et al., 2019), and Relaxation (Bruno and Chaudhuri, 2005) as representatives of this group. On the other hand, AutoAdmin (Chaudhuri and Narasayya, 1997) and Drop (Choenni et al., 1993) serve as baselines for the DROP strategy, where advisors start with an initial set of indexes and iteratively eliminate those deemed inefficient. Learning-based baselines: We adopt two RL-based index advisors as our baselines, namely SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b). While SWIRL includes vast amounts of workloads during training to largely shorten the inference time, BALANCE utilizes transfer RL to increase the generalisability to the dynamic workload scenarios. Prompt-based baselines: The recent work LLMIdxAdvis (Zhao et al., 2025) proposes to recommend indexes by prompting the LLM with prepared demonstrations. Selecting the best demonstration for the input workload and calling the LLM helps LLMIdxAdvis achieve higher efficiency and generalisability. Due to the high computation cost for preparing new demonstrations for our workloads, we directly use LLMIdxAdvis’ prepared demonstrations since they cover all the databases in our experiment.
7.1.4. Evaluation metrics.
We adopt two metrics for our evaluations. The benefit-to-cost ratio metric is first introduced by (Skelley, 2000) and evaluates the absolute workload execution time reduction per unit storage used. The larger the benefit-to-cost ratio, the better an index advisor performs in both workload cost reduction and efficient storage use. The relative workload improvement is defined as the percentage of the workload’s actual execution time decrease before and after the construction of the recommended indexes. We measure all the workload execution time in seconds and storage cost in MegaBytes (MB).
7.1.5. LLM and Training Setting.
We leverage the GPT-4o-0806 (Achiam and et al., 2023) version within the OpenAI API as the default LLM backbone for the agents. Furthermore, we also assess our system’s generalizability across other LLMs, as detailed in Section 8.3. For the index indicator, we train the model on a server with one Nvidia A100 GPU.
@@ 7.1. Experimental Setup
| MAAdivsor | 0.220 | 0.587 | 0.587 | 0.299 | 9.726 | 4.756 | 1.594 | 1.594 | 0.919 | 1.224 | 0.598 | 0.344 | 0.617 | 0.452 | 0.125 | 0.094 |
@@
7.2. Effectiveness and Generalisation Evaluation
@@
As shown in Table 3, we compare our proposed MAAdvisor’s benefit-to-cost ratio with the baselines on four benchmarks. All the methods are evaluated with four different storage budgets, taking 2%, 10%, 20%, and 50% of the database size. Our result reveals several key observations:
@@
(1) MAAdvisor outperforms all the baselines in most of the situations. The benefit-to-cost ratio for our method is the highest in 10 out of the 16 experiment settings, and remains the top tier for the rest 6 settings. Specifically, MAAdvisor doubles the index scores from the SOTA performance given 2%, 10%, and 50% storage budgets in the most complicated DSB benchmark. In addition, similar improvement can be observed in the 10% and 50% budget settings for the real-world benchmark JOB. It demonstrates that MAAdvisor can generalize to a variety of database schemas and workloads while maintaining excellent performance.
@@
| MAAdivsor | 31.85 | 44.31 | 122.21 | 156.82 | 51.00 | 85.75 | 106.71 | 133.21 | 43.23 | 83.67 | 115.38 | 178.82 | 69.22 | 66.18 | 106.62 | 211.56 |
(2) Index regressions happen for the zero-shot heuristic-based methods, for example, Anytime in TPC-H and DSB under 10% budget. Similar results can be observed for Extend in TPC-H and for SWIRL in TPC-DS under 10% budget. Since heuristic-based methods highly rely on the benefit estimations from optimizers, index regressions always happen when the index advisors have a limited storage budget and select the regressive indexes. As an example, DB2Advis recommends indexes for TPC-H that lead to a nearly 50% execution time regression on query template 9 from 14.76 seconds to 26.60 seconds. Among the 13 indexable columns in query 9, only two columns appear in DB2Advis’ index recommendation. In addition, the composite indexes DB2Advis selects ‘I(C supplier.s_nationkey,C supplier.s_suppkey)’ and ‘I(C part.p_type,C part.p_partkey)’ are not efficient for JOIN conditions ‘ps_suppkey = l_suppkey’ and ‘p_partkey = l_partkey’. Unfortunately, the improvements for the indexes on other queries are not significant and therefore result in an overall regression. In contrast, MAAdvisor successfully alleviates the index regression problem while maintaining the zero-shot generalisation ability, with the help of the specifically designed workload representation technique and local revision agent.
(3) Aligned with (Zhou et al., 2024), it is observed that the performance for the learning-based index advisors suffers from higher fluctuation across benchmarks compared to heuristic-based methods. As an example, while BALANCE performs well on TPC-DS and DSB, it fails to beat heuristic-based baselines for the rest of the benchmarks. The main reason is that the training process for the learning-based method is highly variable. The hallucination of LLMs is also a similar issue that may affect the variance of the performance. As an example, LLMIdxAdvis’ performance on TPC-DS 20% budget is among the top tier, but the performance on the other budget settings is low. In contrast, with the help of the hierarchical multi-agent pipeline, our zero-shot MAAdvisor maintains a higher performance across the datasets without training and is robust against hallucination.
@@
(1) MAAdvisor achieves the highest relative workload improvement in 8 out of in total 16 experiment settings. Given smaller storage budgets from 2% to 20%, MAAdvisor tends to recommend indexes that lead to the most workload relative reduction. However, for a larger budget like 50%. Our method prefers indexes with comparable improvement but much lower storage cost. It is also observed that in the experiment settings where our method did not achieve the highest benefit-to-cost ratio, for example, TPC-H 20% budget and JOB 20% budget, MAAdvisor can still recommend indexes that result in the largest workload execution time decrease, with a trade-off of higher storage cost.
(2) Heuristic-based baselines following the DROP heuristic fail to perform well for small storage budgets. The main reason is that the small-storage indexes are generally not as beneficial as large-storage indexes and are removed from the recommendation in early DROP stages. Our MAAdvisor assembles the ADD heuristic that iteratively selects indexes into the recommendation by calling the selection agent. Therefore, MAAdvisor can achieve SOTA performance even for small storage budgets.
(3) The prompt-based method LLMIdxAdvis performs well on TPC-H and JOB benchmarks but has a lower performance on TPC-DS and DSB. As also specified in (Zhao et al., 2025), directly calling LLMs to select indexes fails to excel in these complicated workloads and database schemas. In contrast, MAAdvisor outperforms all the baselines, especially in DSB, with our special design. The workload representation technique efficiently summarizes the information from workloads and database schemas. In addition, the multi-agent pipeline decomposes the problem into sub-steps handled by specialized agents, simplifying the task for each LLM agent to solve.
(4) We find LLMIdxAdvis also struggles to control the storage budget used in TPC-DS and DSB. In TPC-DS, LLMIdxAdvis uses 1356.38 MB under 10% budget, which exceeds the budget, and a lower 1122.97 MB under 20% budget. The main reason is that although the available storage budget is an input for LLM in LLMIdxAdvis, the inputs do not specify the storage cost for selecting new indexes. In contrast, MAAdvisor’s planning and selection of indexes is always storage-aware since the estimated storage cost is recorded in each one of the column candidates.
@@
7.3. Efficiency Evaluation
In Table 4, we record the algorithm inference runtime for all the methods to evaluate their efficiency. Through analyzing the inference runtime of different methods, we have the following key observations:
(1) We can observe that our MAAdvisor’s algorithm runtime outperforms the prompt-based baseline LLMIdxAdvis in most of the settings. The reason is that with the multi-agent pipeline, our method’s only time overheads come from invoking the LLM API and refreshing column candidates, whereas LLMIdxAdvis faces an efficiency bottleneck due to the added cost of identifying the best demonstrations for each selection. In addition, LLMIdxAdvis suffers from high latency during its demonstration construction stage. It is reported that 17 hours and $90 are taken to collect the demonstrations (Zhao et al., 2025). It is also important to note that LLMIdxAdvis’ time cost, given increasing storage budgets, does not change a lot since they call the same demonstration match algorithm. In contrast, our MAAdvisor’s algorithm runtime may increase as the budget increases. It is because our index recommendation is storage-aware. The more budget given, the more local agents are planned by the planning agent. However, MAAdvisor’s time cost under the largest 50% budget is still lower than LLMIdxAdvis.
(2) The learning-based methods have significantly faster inference time compared to the other methods. The reason is that the learning-based methods only need to call the trained models with low latency during inference. However, in a zero-shot situation, the training time of the learned models shall also be considered as part of the algorithm latency. As an example, SWIRL’s training times for the TPC-H, TPC-DS, and JOB datasets are 1226.136, 3373.32, and 3635.00 seconds, respectively. In addition, both learned methods require an initial training set of data collected from the database, which is difficult to acquire in a zero-shot.
(3) Heuristic-based advisors on larger workloads generally suffer from much higher latency. The algorithm runtime for Relaxation on TPC-DS with an 83-query workload is more than 100 times the runtime on the TPC-H benchmark with 19 queries. For a larger workload and more complicated database, MAAdvisor presents a significant efficiency advantage. Although the DB2Advis baseline demonstrates a shorter runtime compared to our MAAdvisor, it trades off with a lower workload performance by using a simpler greedy optimization.
pipeline_example
We further evaluate the efficiency of our MAAdvisor with the LLM baseline LLMIdxAdvis in terms of monetary cost. The average number of input tokens and corresponding API cost for MAAdvisor and LLMIdxAdvis is also recorded in Table 5. It is observed that LLMIdxAdvis’ input lengths are more stable than MAAdvisor across benchmarks. The main reason is that LLMIdxAdvis uses a similar number of demonstrations in all benchmarks. In contrast, MAAdvisor’s input length depends on database complication and the number of column candidates. As an example, MAAdvisor takes 4 times more tokens for TPC-DS with 87 column candidates than TPC-H with 15 column candidates. Considering the complexity of the index recommendation task, MAAdvisor’s additional cost on more complicated workloads is totally acceptable.

### v2

@@ 6.3. Multi-agent Planning and Coordination
Once the high-level and low-level agents are developed, they can be integrated into a hierarchical multi-agent pipeline enabling efficient and effective zero-shot index recommendation. As shown in Algorithm 1, the multi-agent pipeline receives the storage budget and the column candidates as inputs. To balance efficiency with effectiveness, we introduce a hyperparameter , which defines the maximum number of planning sub-steps. At the beginning of each sub-step, the planning agent determines the next action in line 7. If additional indexes can be selected, the selection agent chooses a single-column index from the effectively represented column candidates from line 9 to line 11. When combining existing indexes into a composite index could further improve workload performance, the combination agent in line 13 refines the current recommendation by composing these indexes. Since the selected indexes may contain duplicates, regressions, or low-impact choices, the revision agent ensures effectiveness by removing such cases once called in line 16. After each action, the reflection agent reviews prior decisions and provides guidance for subsequent planning. Through the planning agent’s reasoning in line 22, if the index configuration is deemed optimal, iterations terminate early, and the final selection is output as the index recommendation.
@@
| AMAZe | 0.220 | 0.587 | 0.587 | 0.299 | 9.726 | 4.756 | 1.594 | 1.594 | 0.919 | 1.224 | 0.598 | 0.344 | 0.617 | 0.452 | 0.125 | 0.094 |
7. Experiment
7.1. Experimental Setup
Dataset. We select four different datasets from both hypothetical and real-world settings for our evaluations.
TPC-H: We generate the TPC-H data with a scale of 10 (around 10 GB). The TPC-H benchmark has in total 22 different query templates. In order to simplify the analysis and make a fair comparison of the methods, we follow previous methods (Siddiqui et al., 2022b; Schlosser et al., 2019) to remove the query templates 2, 17, 20 that contribute significantly to the total execution time of the workload or have little potential improvement from indexes. The test workload for TPC-H has in total of 19 queries randomly generated from the 19 query templates left.
TPC-DS: TPC-DS data with a scale of 10 (around 10 GB) is generated. The TPC-DS benchmark has 99 different query templates, and for the same reason in the TPC-H benchmark, we remove the query templates 4, 6, 9, 10, 11, 14, 23, 24, 32, 35, 36, 39, 41, 70, 86, and 95. The test workload for TPC-DS consists of 83 queries generated from the 83 query templates left.
DSB: The DSB benchmark (Ding et al., 2021) is derived from the TPC-DS benchmark, while enhanced with complex data distribution and challenging query templates. DSB has in total 15 templates for single block queries, 22 templates for multi-block queries, and another 15 templates for SPJ queries, especially evaluating techniques with limited capabilities. In our experiments, we only utilized the general-purpose 37 query templates in the first two groups and constructed a workload of 37 queries randomly generated from the templates.
IMDB (JOB workload): Join order benchmark (JOB (Leis et al., 2015)) is an OLAP benchmark querying the IMDB (Maas et al., 2011) dataset and designed to assess the capabilities of a database management system (DBMS) in optimizing and executing complex real-world join queries. The JOB workload has 113 real-world queries.
Experiment Environments. We choose the popular DBMS PostgreSQL 14.2 to evaluate all the queries (Pos, [n.d.]). We follow the official documentaries and websites (TPC, [n.d.]; Leis et al., 2015; dsb, [n.d.]) to create PostgreSQL databases for all the above datasets. All the experiments are conducted on an isolated server with two Intel(R) Xeon(R) Platinum 8469C CPUs featuring 46 cores and 96 threads, along with 460GB of RAM.
Baseline Methods. (1) Heuristic-based baselines: Heuristic-based index advisors can be broadly categorized into two types by their optimization approaches. Those following an ADD strategy incrementally build a set of indexes by selecting candidates one by one. In our experiments, we include Anytime (Chaudhuri and Narasayya, 2020), DB2Advis (Skelley, 2000), Extend (Schlosser et al., 2019), and Relaxation (Bruno and Chaudhuri, 2005) as representatives of this group. On the other hand, AutoAdmin (Chaudhuri and Narasayya, 1997) and Drop (Choenni et al., 1993) serve as baselines for the DROP strategy, where advisors start with an initial set of indexes and iteratively eliminate those deemed inefficient. We reproduce all the heuristic-based baselines based on the evaluation work (Kossmann et al., 2020). (2) Learning-based baselines: We adopt two RL-based index advisors as our baselines, namely SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b). While SWIRL includes vast amounts of workloads during training to largely shorten the inference time, BALANCE utilizes transfer RL to increase the generalisability to the dynamic workload scenarios. (3) Prompt-based baselines: The recent work LLMIdxAdvis (Zhao et al., 2025) proposes to recommend indexes by prompting the LLM with prepared demonstrations. Selecting the best demonstration for the input workload and calling the LLM helps LLMIdxAdvis achieve higher efficiency and generalisability. Due to the high computation cost for preparing new demonstrations for our workloads, we directly use LLMIdxAdvis’ prepared demonstrations since they cover all the databases in our experiment.
Evaluation metrics. We adopt two metrics for our evaluations. The benefit-to-cost ratio metric is first introduced by (Skelley, 2000) and evaluates the absolute workload execution time reduction per unit storage used. The larger the benefit-to-cost ratio, the better an index advisor performs in both workload cost reduction and efficient storage use. The relative workload improvement is defined as the percentage of the workload’s actual execution time decrease before and after the construction of the recommended indexes. We measure all the workload execution time in seconds and storage cost in MegaBytes (MB).
LLM and Training Setting. We leverage the GPT-4o-0806 (Achiam and et al., 2023) version within the OpenAI API as the default LLM backbone for the agents. Furthermore, we also assess our system’s generalizability across other LLMs, as detailed in Section 7.5. For the index indicator, we train the model on a server with one Nvidia A100 GPU.
@@ 7.1. Experimental Setup
7.2. Generalization and Effectiveness Evaluation
@@
As shown in Table 3, we compare our proposed AMAZe’s benefit-to-cost ratio with the baselines on four benchmarks. All the methods are evaluated with four different storage budgets, taking 2%, 10%, 20%, and 50% of the database size. Our results reveal several key observations:
@@
(1) AMAZe outperforms all the baselines in most of the situations. The benefit-to-cost ratio for our method is the highest in 10 out of the 16 experimental settings, and remains the top tier for the rest 6 settings, especially compared to the heuristic-based ßand prompt-based methods, which are also zero-shot. Specifically, AMAZe doubles the ratios from the SOTA performance given 2%, 10%, and 50% storage budgets in the most complicated DSB benchmark. Similar improvement can be observed in the 10% and 50% budget settings for the real-world benchmark JOB. This demonstrates AMAZe’s strong generalization capability to a variety of database schemas and workloads in zero-shot settings.
(2) AMAZe successfully alleviates the index regression problem, which commonly occurs in the zero-shot heuristic-based methods. For example, Anytime suffers from the index regression in TPC-H and DSB under 10% budget, and similar problems can be observed for Extend in TPC-H and for SWIRL in TPC-DS under 10% budget. Our detailed analysis reveals that heuristic-based methods highly rely on the benefit estimations from optimizers, index regressions always happen when the index advisors have a limited storage budget and select the regressive indexes. For example, DB2Advis recommends indexes for TPC-H that lead to a nearly 50% execution time regression on query template 9 from 14.76 seconds to 26.60 seconds. Among the 13 indexable columns in query 9, only two columns appear in DB2Advis’ index recommendation. In addition, the composite indexes DB2Advis selects ‘I(C supplier.s_nationkey,C supplier.s_suppkey)’ and ‘I(C part.p_type,C part.p_partkey)’ are not efficient for JOIN conditions ‘ps_suppkey = l_suppkey’ and ‘p_partkey = l_partkey’. Unfortunately, the improvements for the indexes on other queries are not significant and therefore result in an overall regression. In contrast, our AMAZe mitigates the problem with the help of the specifically designed workload representation technique and low-level revision agent.
(3) AMAZe mitigates the training instability or hallucination problem faced by learning-based / prompt-based methods. Our results show that the performance for the learning-based index advisors suffers from higher fluctuation across benchmarks. As an example, while BALANCE performs well on TPC-DS and DSB, it fails to beat heuristic-based baselines in the rest of the benchmarks due to the highly variable training process (Zhou et al., 2024). Similar variation happens in the prompt-based methods due to the hallucination of LLMs. As an example, LLMIdxAdvis’ performance on TPC-DS 20% budget is among the top tier, but the performance on the other budget settings is low. In contrast, our hierarchical multi-agent pipeline verifies and corrects the potential hallucinations during the index recommendation, effectively mitigating the problem.
(1) AMAZe achieves the highest relative workload improvement in 8 out of in total 16 experiment settings. Given smaller storage budgets from 2% to 20%, AMAZe tends to recommend indexes that lead to the most workload relative reduction. However, for a larger budget like 50%. Our method prefers indexes with comparable improvement but much lower storage cost. It is also observed that in the experiment settings where our method did not achieve the highest benefit-to-cost ratio, for example, TPC-H 20% budget and JOB 20% budget, AMAZe can still recommend indexes that result in the largest workload execution time decrease, with a trade-off of higher storage cost.
(2) AMAZe achieves SOTA performance even for small storage budgets. In contrast, heuristic-based baselines following the DROP heuristic fail to perform well for small storage budgets. The main reason is that the small-storage indexes are generally not as beneficial as large-storage indexes and are removed from the recommendation in early DROP stages. Our AMAZe assembles a hybrid of both ADD and DROP heuristics that iteratively selects and revises the index recommendation by multi-agent coordinations.
(3) AMAZe with the multi-agent design outperforms the prompt-based baseline, especially in the most complicated DSB. The prompt-based LLMIdxAdvis performs well on TPC-H and JOB benchmarks but has a lower performance on complicated workloads and database schemas, like TPC-DS and DSB. AMAZe’s better performance comes from the efficient workload representation technique which summarizes the information from workloads and database schemas. In addition, the multi-agent pipeline decomposes the problem and simplifies the task for each LLM agent.
(4) AMAZe better controls the storage budget usage during the recommendation. It is observed that LLMIdxAdvis struggles to control the storage budget used in TPC-DS and DSB. As an example in TPC-DS, LLMIdxAdvis uses 1356.38 MB under 10% budget, which exceeds the budget, and a lower 1122.97 MB under 20% budget. The main reason is that although the available storage budget is an input in LLMIdxAdvis, LLMs are not given the storage cost when selecting new indexes. In contrast, AMAZe’s planning and selection of indexes is always storage-aware since the estimated storage cost is recorded in each one of the column candidates.
@@
| AMAZe | 31.85 | 44.31 | 122.21 | 156.82 | 51.00 | 85.75 | 106.71 | 133.21 | 43.23 | 83.67 | 115.38 | 178.82 | 69.22 | 66.18 | 106.62 | 211.56 |
7.3. Efficiency Evaluation
Table 4 reports the algorithm inference runtime for all the methods to evaluate their efficiency. Through analyzing the inference runtime of different methods, we make the following observations:
(1) AMAZe’s runtime outperforms the prompt-based baseline LLMIdxAdvis in most of the settings. The reason is that with the multi-agent pipeline, our method’s main computational overhead comes from invoking the LLM API and updating column candidates, whereas LLMIdxAdvis faces an efficiency bottleneck due to the added cost of identifying the best demonstrations for each selection. In addition, LLMIdxAdvis suffers from high latency during its demonstration construction stage. 17 hours and $90 are reported to be taken to collect the demonstrations (Zhao et al., 2025). It is also note-worthy that LLMIdxAdvis’ time cost, given increasing storage budgets, does not change a lot since they call the same demonstration match algorithm. In contrast, AMAZe’s runtime may increase as the budget increases. It is because our index recommendation is storage-aware. The more budget given, the more low-level agents are planned by the planning agent. However, AMAZe’s time cost under the largest 50% budget is still lower than LLMIdxAdvis.
(2) The learning-based methods have significantly faster inference time compared to the other methods. The reason is that the learning-based methods only need to call the trained models with low latency during inference. However, in a zero-shot situation, the training time of the learned models shall also be considered as part of the algorithm latency. As an example, SWIRL’s training times for the TPC-H, TPC-DS, and JOB datasets are 1226.136, 3373.32, and 3635.00 seconds, respectively. In addition, both learned methods require an initial training set of data collected from the database, which is difficult to acquire in a zero-shot setting.
(3) AMAZe presents a significant efficiency advantage for larger workloads and more complicated database schemas. Heuristic-based advisors on larger workloads generally suffer from much higher latency. As an example, the algorithm runtime for Relaxation on TPC-DS with 83 queries’ workload is more than 100 times the runtime on the TPC-H benchmark with 19 queries. Although the DB2Advis baseline shows a shorter runtime compared to our AMAZe, it trades off with a lower workload performance by using a simpler greedy optimization.
We further evaluate the performance of AMAZe with the LLM baseline LLMIdxAdvis in terms of monetary cost. The average number of input tokens and the corresponding API cost for AMAZe and LLMIdxAdvis are recorded in Table 5. It is observed that LLMIdxAdvis’ input lengths are more stable across benchmarks. The main reason is that LLMIdxAdvis uses a similar number of demonstrations in all benchmarks. In contrast, AMAZe’s input length depends on database complication and the number of column candidates. As an example, AMAZe takes 4 times more tokens for TPC-DS with 87 column candidates than TPC-H with 15 column candidates. Considering the complexity of index recommendation task, AMAZe’s higher cost on more complicated workloads is acceptable.
@@ 7.3. Efficiency Evaluation
pipeline_example